# New theoretical Fe II templates for bright quasars

### v1

@@ 3.8.1 HE 0413-4031
This quasar has been observed with the Southern African Large Telescope (SALT) for 13 years with the aim of performing the reverberation mapping of the Mg II line and the UV Fe II (Zajaček et al., 2020; Prince et al., 2023; Zajaček et al., 2023b). The obtained spectra cover a narrow range around the Mg II line but the spectral shape did not show any strong variability apart from the normalization of the spectrum. We thus constructed the mean spectrum of this quasar by combining all 32 available data sets for a better S/N ratio. The shape of the Mg II line in this source is particularly simple, well fitted as a single kinematic component of the Lorentzian shape.
@@
We fit all the parameters together (slope and normalization of the power law, normalization and the broadening of the Fe II template, the position, normalization and FWHM of the Mg II) since in case of such a short data range we do not have any possibility to find parts of the spectrum unaffected by Fe II and fit first the underlying power law.
@@
The narrow spectral range makes the fits not strongly constraining the Fe II templates but the fits are very satisfactory. The results of the spectral fitting are shown in Figure 13. The best solution corresponds to , shown in the middle panel of Figure 13. The best fit FWHM of the Fe II (5400 km s) is broader than the FWHM of Mg II (4280 km s) but it comes with a large error. Higher turbulent velocity (100 km s) gives fits with somewhat higher reduced , and pushes the FWHM of Fe II toward even higher values, at the same time increasing the Fe II intensity. This results from considerable degeneracy between the Fe II normalization and the underlying power law. A lower value of the turbulent velocity (20 km s) gives narrower Fe II lines but the quality of the fit is much worse.
@@
For comparison, we show two other fits, for higher ( (cm s) = 19) and lower ( (cm s) = 18) ionization fluxes in the upper and lower panels of Figure 13, respectively. The fit quality is good in all these cases ( differences ), and the value of the EW of the Mg II line is similar (31.2 Å, 30.2 Å and 28.4 Å), but slowly rising towards larger ionization fluxes. We calculated also EW(Fe II) but only in the fitted range. EW(Fe II) is more affected by the change of the ionizing flux (the corresponding values are 37.9 Å, 31.2 Å and 15.9 Å), and all these values are larger than achieved with the use of different templates by Zajaček et al. (2023b). Such a large change is related to a corresponding change in the slope and normalization of the underlying power law.
@@
The chosen fit density ( (cm) = 12) is consistent with the best fit theoretical template d12-m20-20-5.dat selected from the collection of Bruhweiler & Verner (2008) in our previous papers, including Prince et al. (2023). However, the ionization flux is now lower by more than two orders of magnitude.
@@
We compared the requested value for the ionizing flux from Fe II modelling with the expected ionizing flux based on SED of this object and the Mg II time delay of 302.6 days for Mg II (Zajaček et al., 2020) but the obtained ionization flux from the SED best fit presented in Figure 9 of Zajaček et al. (2020) gave only , since the spectrum appeared to decrease rapidly in the UV band.
The narrow available spectral range on one hand guarantees that a range of templates fits well the spectrum but it opens a question of how reliable is such a decomposition.
3.8.2 SDSS RM102
As a second example, we selected the quasar SDSS J141352.99+523444.2 (SDSS RM102) from the SDSS Reverberation Mapping (SDSS-RM) catalogue (Shen et al., 2019) considering a high signal-to-noise, moderate Fe II emitter (), and a reliable time lag for H or Mg II (maximum Pearson correlation coefficient ). The spectrum of the source covers a broad wavelength range (2000-5500 Å). This allowed us to cover both the rest-frame optical Fe II as well as UV Fe II and to test the feasibility of using the same Fe II template to fit the entire UV-to-optical spectral range. The size of its BLR region has been estimated by different methods, in this paper we took the one estimated by the Interpolated Cross-Correlation Function (ICCF), and days in the observed-frame (Shen et al., 2023). Using the SDSS Science Archive Server (SAS) from the DR16333https://dr16.sdss.org/optical/spectrum/search, we found 70 spectroscopic epochs observed during four years. All the spectra were combined in a single one to obtain a high signal-to-noise spectrum, S/N=30 and 35 at 3000Å and 4000Å, respectively. Due to the presence of spikes around 5100Å, it was not possible to estimate the S/N at this wavelength.
In the spectral fitting, we considered all the emission lines reported by Vanden Berk et al. (2001) between 2000 and 5500 Å, 31 emission lines in total. We also included a model of the Balmer Continuum with an optical depth of 1 plus the High-Order Balmer lines (Bernal et al in prep.). The continuum was fitted with a single power law. To decrease the number of free parameters, we fixed the same FWHM for each family of emission lines: Balmer lines, Helium transitions, forbidden lines (narrow emission lines), and coronal lines. Flux intensities, FWHM, and shift sum a total of 63 free parameters. We applied Gaussian smearing to the UV and optical Fe II template and a single flux intensity value was taken for the UV and optical wavelengths. In the first attempt, we noticed that it was not possible to fit the UV and optical Fe II contribution with the same flux intensity. Thus, we performed a careful fitting of all the other spectral features and what remained can be considered as the Fe II component. Due to the overlapping of the emission lines and the Fe II pseudocontinuum in some ranges, the flux intensity of the emission line would be underestimated. So, we considered the semiempirical UV/optical Fe II templates implemented in the fitting code PyQSOFit (Guo et al., 2018) to constrain the intensity of the emission lines. As we previously noticed, even with these empirical Fe II templates, a different flux intensity value is needed to fit the Fe II contribution. The flux intensity factor has a difference of 3.5 between the UV and the optical Fe II contribution. The final spectral fitting is shown in Figure 14. The potential Fe II contribution is shown at a zero level. As a reference, we also plotted the UV (dotted line) and optical (dashed-line) Fe II templates implemented in PyQSOFit. There is good agreement in the optical Fe II, while slight variations are seen in the UV part.
The new spectral fitting indicates a =1.24, which is approximately twice the value reported by Shen et al. (2019). On the other hand, this source shows a moderate Eddington ratio =0.329 (Shen et al., 2019). The weak contribution of the doublet [OIII] suggests that RM102 has a moderate Eddington ratio and is close to the xA sources (Marziani, 2023).
To locate the source in the plane, we measured three ratios: red to blue wings of UV Fe II, red to blue wings of optical Fe II, and UV to optical Fe II. The first two ratios were measured in the same wavelength range as used to create the general maps in Figure 2 and Figure 5. The last ratio had to be modified since the data cover the wavelength range only up to 2000 Å. The measured ratios are given in Table 4. Looking at the plots in Section 3 above we could not easily identify the most likely parameters of the successful template. We thus created a new map, shown in Figure 15, where all the three ratios given in Table 4 are shown as contour plots, each for all 6 turbulent velocities.
wavelength range 1 | wavelength range 2 | ratio 1 to 2 |
2800-3090 Å | 2650-2800 Å | 1.939 |
5150-5350 Å | 4434-4684 Å | 0.713 |
2000-3090 Å | 4000-5350 Å | 2.151 |
We see that the three families of lines in Figure 15 never cross. This means that there is no single template in our collection which could fit the data for the quasar RM102. The blue lines representing the ratio of red to blue wings of optical Fe II occupy a narrow range of high ionization flux values. This reflects the fact that the measured ratio is low and the corresponding solutions are located in the upper part of the general diagram (see Figure 5). On the other hand, the UV wings ratio in general forms a two-family solution for low values (see Figure 2), but for high values as the one measured in our source there is only an island located at low ionization flux and low density. The global UV to optical ratio (although measured in a slightly different wavelength range than the generic Figure 1) also points toward low density and low ionization flux. The results for densities of the order of cm and moderate ionization fluxes of the order of cm s are similar to the parameter range concluded for HE 0413-4031, but the optical wing ratio remains unexplained.
@@ 3.8.2 SDSS RM102
We estimated the ionization flux from this source at the distance corresponding to the BLR position as given by the H line delay. However, the SDSS spectrum does not cover the spectral range above 1 Rydberg, needed for this purpose. We thus extrapolated the spectrum of RM102 using the AGN SED template from CLOUDY. The observed spectrum is only marginally bluer in the optical band. Matching the two spectra at UV part we obtained the ionization flux of cm s. We also tested another way of extension, using the HST composite spectrum of Zheng et al. (1997) which covers the rest frame range from 350 Å up to 2200 Å, and they give the broken power law fit. Matching this broken power law to the shortest wavelengths with our spectrum we obtained the ionization flux of cm s. The two values are comparable, they eventually favour the value implied by the red-to-blue optical Fe II ratio, although at (rather likely) higher densities than cm they are not in full agreement with any of the ratios. However, some composites predict a stronger decrease of the quasar far-UV flux towards shorter wavelengths (see Cai & Wang, 2023, and the references there).
@@
Overall, Figure 15 implies that a single Fe II template is unlikely to fit well the whole spectrum. In particular, the mismatch in the values of the ionization flux is by a factor of 100. The metallicity is not affecting the ratios strongly enough. We thus set the density at the likely value of cm, and we took two approaches to set the ionizing flux. In the first case, we allowed for a discontinuity in the ionizing flux at 4000 Å, where the Fe II emission is very low, and we used low ionizing flux in UV band () and high ionizing flux () in the optical band. Separate fitting of the Fe II and optical is usually done when modelling a specific source. The result of this approach is shown in Figure 16. In the second approach, we allowed for a linear change of the ionizing flux with the wavelength in the log-log scale. This result is shown in the same plot (magenta line). None of these attempts are fully satisfactory. However, the discontinuous model seems to work much better. Still, in both cases, we face a considerable mismatch between the model and the data in the region 3000 - 3500 Å. Semi-empirical models do include considerable emission there (see Figure 14, dotted line). The spike close to 2000 Å is most likely a Fe III emission. We did not try a more complex approach in the form of the Locally Optimized Cloud (LOC) model of Baldwin et al. (1995) since we do not include Fe III emission which may be important for exact data fitting. We address this issue and additional likely effects in the discussion.
@@ 4 Discussion
We used the most recent version of the code CLOUDY C23.00 to construct the grid of theoretical Fe II templates covering the broad UV to NIR range from 1000 Å to 10000 Å. The templates represent single-zone emitters, parametrized by the local density and ionization flux. We consider a range of turbulent velocities and for a fixed turbulent velocity of 20 km we calculated the results for different metallicities. For the other turbulent velocity the considered metallicity is solar.
@@
The advantage of the theoretical templates is that they predict the Fe II emission underlying the strong emission lines while purely observational templates, modelling first the lines, force low Fe II emission at the position of lines like H or Mg II. In addition, the construction of the observational templates is thus most successful for the objects with extremely narrow emission lines like I Zw 1 (Vestergaard & Wilkes, 2001) or Mrk 493 (Park et al., 2022). Line widths of these spectra imply high Eddington ratio, strong Fe II emission and the incident SED shape not necessarily representative for the majority of quasars.
@@
The availability of the Fe II templates in the full spectral range will address better the issue of the production of the Fe II in the BLR. In the case of the observational data covering a narrow spectral range, the templates allow us to fit the data and obtain an estimate of the density and the ionization flux in the BLR, as we showed in Section 3.8.1. However, our attempt to use these templates for a quasar spectrum covering both optical and UV ranges was not successful (see Section 3.8.2). The problem seems generic since the selected quasar is rather a typical object, and the issue of the UV to optical ratio in the models and the data has been present in the literature for many years.
Already the early papers asked the question of whether the Fe II emission actually forms as a result of the irradiation by the central source or due to mechanical heating. For example, Joly (1981) in her paper argued that either mechanical heating is the dominating channel of production, or the clouds must be very optically thick, in which case the irradiation flux is thermalized and the mechanism is similar to the mechanical heating. However, those early models were not suitable for column densities larger than cm.
Verner et al. (2003) calculated the predicted ratio of the UV to optical Fe II emission from their model of the Fe II atomic transitions. The adopted ranges were 2000–3000 Å and 4000–6000 Å range, the range is comparable although not identical to the one studied in Section 3.8.2. The UV/optical ratio obtained in this paper is similar to what we measure from RM102 in the case of collisional heating, but radiative heating gives much higher values, 8 - 11, depending on the number of transitions. These values were obtained under a turbulent velocity of 1 km s. They noted that Fe II optical emission flux is more sensitive to abundance than the Fe II (UV) band. Conversely, the Fe II (UV) band is more sensitive to microturbulence than the Fe II optical band.
Baldwin et al. (2004) investigated the Fe II emission in the UV (2200-2800 Å) and optical (4450-4750 and 5080-5460 Å) ranges. They obtained the observed Fe II UV/optical ratio of 1-2, which is significantly lower than the model-predicted ratio of 40-50. Using the LOC model, they found that the observed UV Fe II emission requires a high microturbulence velocity ( 100 km/s). They examined collisional heating as an alternate excitation mechanism, however only the Fe II lines were visible in the UV range, which has the disadvantage of adding another specialised region to the already complicated structure of the inner region of AGN. They noticed that other complex ions, such as Ni II, might have created similar patterns despite their lower abundance, but their atomic data was not available. As a result, they favoured a microturbulence-based model for explaining UV Fe II emissions and concluded that the optical Fe II emission originates from an entirely different region compared to the UV Fe II emission.
The problem of matching the theoretical templates over the full optical to UV spectral range is not likely due to the use of a single-zone approximation, particularly as some of the papers mentioned above used a multi-zone approach. The standard LOC model can in principle represent both the broad range of radii as well as the broad range of densities but may not be actually appropriate to solve the physical problem.
The radially very extended emission is unlikely. Kovačević-Dojčinović & Popović (2015) argued that there is a kinematic correlation between the UV and the optical Fe II emissions, although they reported considerable positive shift of the UV Fe II with respect to the systemic redshift for the majority of their sources (their Fig. 10). The radius-luminosity relation discussed by Zajaček et al. (2023b) for both spectral ranges looks roughly similar, so the two regions cannot be considerably shifted. Furthermore, the possibility of measuring the Fe II delay in the optical and UV bands contradicts the major collisional origin of the emission, but some contributions cannot be ruled out. The atomic content of the recent version C23.00 of CLOUDY used in this paper does not seem questionable, and apparently, the systematic improvement in the data did not solve the long-standing issue of UV and optical discrepancy. The metallicity dependence also does not seem to be strong enough to account for the issue.
There are thus two remaining issues which can solve the problem but they were not modelled in this paper since our primary goal was to provide a set of theoretical templates for further use.
The first weakness in the simulation is the use of a constant density assumption. Such an approach is widely employed, yet it is not physically warranted. The higher temperature of the illuminated face of the cloud and the considerably lower temperature of the dark side would create an enormous pressure gradient which could disrupt the cloud during the dynamical timescale. Instead, the illuminated cloud must be in pressure equilibrium which also takes into account the radiation pressure (Różańska et al., 2006; Baskin et al., 2014). Overall, the bright face has a much lower density, and the density is rising towards the dark side. Clouds in pressure equilibrium were considered in several papers. Some of these computations were done in the context of the warm absorber (Różańska et al., 2006; Adhikari et al., 2015, 2019) but this medium may be the same as BLR. The change of the density and temperature is not like a power law but rather like a rapid transition (see e.g. Fig. 1 of Baskin et al. 2014, or Fig. 3 of Adhikari et al. 2019), so the LOC model would not reproduce it well, and eventually, a discontinuous jump in density and ionization parameter might be a better approximation. Computations of clouds under constant pressure are possible within CLOUDY environment, and this could be possibly done in the future.
@@
The second weakness is the assumption of isotropic emission. If the clouds are few and randomly distributed around the central source, then we will see clouds at all angles with respect to their orientation towards the source. The entire contribution summed up is well represented by the isotropic emission. If cloud distribution is flattened, but our line of sight is not passing through the cloud region the dark sides may contribute a little stronger to the total emission in terms of the effective surface but most of the cloud emission comes from the bright sides. Such expectations were given, for example, by Czerny & Hryniewicz (2011) in the context of the time delay. However, the bright sides contribute more to the isotropic part, therefore overexposure of the cloud’s bright sides may not have a significant impact on Fe II emission. However, Ferland et al. (2009) suggested that we predominantly see the dark side of the clouds. They connected this to the Fe II emission in the optical band showing predominantly the redshift (Hu et al., 2008) and thus implying the inflow. Also in Panda et al. (2018), in Section 4.3, they discussed the issue that the optical Fe II is produced predominantly in the dark side of a cloud, and the dominance of the dark side helped to reproduce ratio without metallicity enhancement. However, the requested geometrical setup is not simple since the line of sight does not cross the BLR. It may require a highly ionized absorber, between the black hole and the BLR, almost ‘invisibly’ shielding the well-irradiated bright sides of clouds.

### v2

@@ 3.10.1 HE 0413-4031
The observed spectrum of HE 0413-4031 covers a narrow range around the Mg II line but the spectral shape did not show any strong variability apart from the normalization of the spectrum. We thus constructed the mean spectrum of this quasar by combining all 32 available data sets for a better S/N ratio. The shape of the Mg II line in this source is particularly simple, well fitted as a single kinematic component of the Lorentzian shape.
@@
We fit all the parameters together (slope and normalization of the power law, normalization and the broadening of the Fe II template, the position, normalization, and FWHM of the Mg II) since in case of such a short data range we do not have any possibility to find parts of the spectrum unaffected by Fe II and fit first the underlying power law.
@@
The narrow spectral range ( 2700 - 2900 Å), centered around the broad Mg II line, makes the fits not strongly constraining the Fe II templates but the fits are very satisfactory. The results of the spectral fitting are shown in Figure 19. The best solution corresponds to , shown in the middle panel of Figure 19. The best fit FWHM of the Fe II (5400 km s-1) is broader than the FWHM of Mg II (4280 km s-1) but it comes with a large error. Higher microturbulence velocity (100 km s-1) gives fits with somewhat higher reduced , and pushes the FWHM of Fe II toward even higher values, at the same time increasing the Fe II intensity. This results from considerable degeneracy between the Fe II normalization and the underlying power law. A lower value of the microturbulence velocity (20 km s-1) gives narrower Fe II lines but the quality of the fit is much worse.
@@
For comparison, we show in Figure 19 two other fits, for higher ( (cm-2 s-1) = 19) and lower ( (cm-2 s-1) = 18) ionizing photon fluxes in the upper and lower panels of Figure 19, respectively. The fit quality is good in all these cases, the value of measuring the fit quality was higher only by for the upper and the lower panel than for the middle panel), and the value of the EW of the Mg II line is similar (31.2 Å, 30.2 Å and 28.4 Å), but slowly rising towards larger ionizing photon fluxes. We also calculated EW(Fe II) but only in the fitted range. EW(Fe II) is more affected by the change of the ionizing flux (the corresponding values are 37.9 Å, 31.2 Å , and 15.9 Å), and all these values are larger than achieved with the use of different templates by Zajaček et al. (2024b). Such a large change is related to a corresponding change in the slope and normalization of the underlying power law.
@@
The chosen fit density ( (cm-3) = 12) is consistent with the best fit theoretical template d12-m20-20-5.dat selected from the collection of Bruhweiler & Verner (2008) in our previous papers, including Prince et al. (2023). However, the ionizing photon flux is now lower by more than two orders of magnitude than in the template; (cm-2 s-1) = 18 in the model versus (cm-2 s-1) = 20.5 in the template.
@@
We compared the requested value for the ionizing flux from Fe II modeling with the expected ionizing flux based on the SED of this object and the Mg II time delay of 302.6 days for Mg II (Zajaček et al., 2020). The obtained ionizing photon flux from the SED best fit presented in Figure 9 of Zajaček et al. (2020) gave , although the spectrum appears to decrease rapidly in the far UV band. This ionizing flux is higher by two orders of magnitude than implied by the current fits, and consistent with the old Bruhweiler & Verner (2008) template used in Zajaček et al. (2020) and Prince et al. (2023). The narrow available spectral range on one hand guarantees that a range of templates fits well the spectrum but it opens a question of how reliable is such a decomposition.
3.10.2 RM 102
@@
We estimated the ionizing photon flux from this source at the distance corresponding to the BLR position as given by the H line delay. However, the SDSS spectrum does not cover the spectral range above 1 Rydberg, needed for this purpose. We thus extrapolated the spectrum of RM102 using the AGN SED template from CLOUDY. The observed spectrum is only marginally bluer in the optical band. Matching the two spectra at the UV part we obtained the ionizing photon flux of cm-2 s-1. We also tested another way of extension, using the HST composite spectrum of Zheng et al. (1997) which covers the rest frame range from 350 Å up to 2200 Å, and they give the broken power law fit. Matching this broken power law to the shortest wavelengths with our spectrum we obtained the ionizing photon flux of cm-2 s-1. The two values are comparable. However, some composites predict a stronger decrease of the quasar far-UV flux towards shorter wavelengths (see Cai & Wang, 2023, and the references therein).
@@
To compare our model-predicted Fe II spectra with the observed spectra of RM 102, we plotted sample spectra from different models and for microturbulence velocities 50 and 100 km/s in Figure 20. These spectra include a covering factor of 30% and are convolved with a Gaussian function having FWHM = 4000 km s-1. Figure 20 implies that a single Fe II template is unlikely to fit well the whole spectrum. In the UV part, the model-predicted spectra provide a somewhat better match to the observed spectra but in the optical region, the model-predicted Fe II emission is much weaker compared to the observed Fe II. The metallicity is not affecting the ratios strongly enough. We thus set the density at the likely value of cm-3, and we took a different approach to fit the observed spectra. We allowed for a discontinuity in the ionizing flux at 4000 Å, where the Fe II emission is very low. We used a high ionizing flux () in the UV band and a low ionizing flux () in the optical band. Separate fitting of the UV and optical Fe II is usually done when modeling a specific source. We also chose different covering factors in UV (30%) and optical (50%) regions. The need for a higher covering factor in the optical part has been suggested in several studies (Gaskell, 2009; Ferland et al., 2009). The result of this approach is shown in Figure 21. None of these attempts are fully satisfactory. However, the discontinuous model seems to work much better. Still, in both cases, we face a considerable mismatch between the model and the data in the region 3000 - 3500 Å. Semi-empirical models do include considerable emission there (see bottom panel of Figure 1, dotted line). The spike close to 2000 Å in the data is most likely a calibration issue.
@@ 4 Discussion
We used the most recent version of the code CLOUDY C23.00 to construct the grid of theoretical Fe II templates covering the broad UV to NIR range from 1000 Å to 10000 Å. The templates represent single-zone emitters, parametrized by the local density and ionizing photon flux. We consider a range of microturbulence velocities and for a fixed microturbulence velocity of 100 km s-1 we calculated the results for different metallicities. For the other microturbulence velocity the considered metallicity is solar.
@@
The advantage of the theoretical templates is that they predict the Fe II emission underlying the strong emission lines while purely observational templates, modeling first the lines, force low Fe II emission at the position of lines like H or Mg II. In addition, the construction of the observational templates is thus most successful for the objects with extremely narrow emission lines like I Zw 1 (Vestergaard & Wilkes, 2001) or Mrk 493 (Park et al., 2022). Line widths of these spectra imply a high Eddington ratio, strong Fe II emission, and the incident SED shape not necessarily representative of the majority of quasars.
@@
The availability of the Fe II templates in the full spectral range will address better the issue of the production of the Fe II in the BLR. In the case of the observational data covering a narrow spectral range, the templates allow us to fit the data and obtain an estimate of the density and the ionizing photon flux in the BLR, as we showed in Section 3.10.1. However, our attempt to use these templates for a quasar spectrum covering both optical and UV ranges was unsuccessful (see Section 3.10.2). The problem seems generic since the selected quasar is rather a typical object, and the issue of the UV to optical ratio in the models and the data has been present in the literature for many years.
Our template fluxes are properly normalized so we can check the equivalent width of the Fe II emission in various bands corresponding to 100 % covering factor of the central source by the BLR region. Overall, the Fe II emission strength of the irradiated faces of the clouds is large enough to represent the actual emission when the realistic covering factor is included. However, in most bands, it requires relatively high microturbulent velocity, close to 100 km s-1. Such high microturbulence velocity was discussed in some previous works (e.g. Baldwin et al., 2004; Panda et al., 2018, 2019a; Sarkar et al., 2021). A microturbulence of this magnitude must reflect the dispersion of kinematic velocity rather than the thermal motion of the gas. This is not unexpected in the case of a turbulent medium. The medium, most likely formed as a clumpy wind from the disk (e.g. Murray et al., 1995; Czerny & Hryniewicz, 2011; Elvis, 2017) shows systematic Keplerian motion of a few thousand km s-1 but clump formation or failed wind approach can introduce the velocity field being a fraction of this Keplerian speed, and affect the transparency of the medium. Thus the medium represented by a combination of directly irradiated and not irradiated clouds in roughly equal proportion can well represent the BLR. Additionally, the magnetic field likely present in the BLR can lead to the development of internal motions at a level of even a few hundred km s-1 (Krause et al., 2012).
@@
We also studied separately the dark (not directly irradiated) cloud sides since some inflow models and special geometries may suggest that we mostly see the dark sides of the clouds (e.g. Ferland et al., 2009; Wang et al., 2017). If clouds are few and randomly distributed around the central source, then we will see clouds at all angles to their orientation toward the source. The entire contribution summed up is well represented by the isotropic emission. If cloud distribution is flattened, but our line of sight is not passing through the cloud region the shielded sides may contribute a little stronger to the total emission in terms of the effective surface but most of the cloud emission comes from the bright sides. Such expectations were given, for example, by Czerny & Hryniewicz (2011) in the context of the time delay. However, the bright sides contribute more to the isotropic part, therefore overexposure of the cloud’s bright sides may not have a significant impact on Fe II emission. The dominance of the shielded material changes the picture considerably. Ferland et al. (2009) connected this effect to the Fe II emission in the optical band showing predominantly the redshift (Hu et al., 2008) and thus implying the inflow. Also in Panda et al. (2018), in Section 4.3, they discussed the issue that the optical Fe II is produced predominantly in the unilluminated side of a cloud, and the dominance of the unilluminated side helped to reproduce ratio without metallicity enhancement. The requested geometrical setup is not simple since the line of sight does not cross the BLR. It may require a highly ionized absorber, between the black hole and the BLR, almost ‘invisibly’ shielding the well-irradiated bright sides of clouds. However, we show now that if dark sides statistically dominate, the strength of Fe II emission is too low to reproduce the overall Fe II emission in typical quasars.
Moreover, when confronting the templates with the broadband optical/UV data with the test object RM 102 we could not well represent the ratios of the UV-to-optical band, despite the use of the newest atomic data. The problem was already discussed by Joly (1981); Verner et al. (2003); Baldwin et al. (2004). Baldwin et al. (1995) used the Locally Optimized Cloud (LOC) model which allowed for a combination of clouds with a range of densities and distances to solve the issue and concluded that UV and optical emission come from two separate regions of different properties. But the radially very extended emission is unlikely. Kovačević-Dojčinović & Popović (2015) argued that there is a kinematic correlation between the UV and the optical Fe II emissions, although they reported considerable positive shift of the UV Fe II for the systemic redshift for the majority of their sources (their Fig. 10). The radius-luminosity relation discussed by Zajaček et al. (2024b) for both spectral ranges looks roughly similar, so the two regions cannot be considerably shifted. The problem of matching the theoretical templates over the full optical to UV spectral range is not likely due to the use of a single-zone approximation, particularly as some of the papers mentioned above used a multi-zone approach.
However, we should note that there may also be a problem on the observational side. Despite the selection of the object with a high S/N ratio, the determination of the strength of Fe II emission is not perfect, particularly in the optical band where we also have some starlight contribution. We do not think it is the main source of the problem, but certainly, future applications of our new templates using the LOC model and more objects can shed light on the issue.
There are thus two remaining issues that can solve the problem and could be done in principle but they were not modeled in this paper since our primary goal was to provide a set of theoretical templates for further use. We do not discuss below the issues which are known but are far beyond the modeling possibilities. These are the dynamical aspects of cloud evolution: formation and destruction, including the collisions with the accretion disk. For a short summary and the references, see Appendix in Müller et al. (2022).
The first weakness in the simulation is the use of a constant density assumption. Such an approach is widely employed, yet it is not physically warranted. The higher temperature of the illuminated face of the cloud and the considerably lower temperature of the shielded side would create an enormous pressure gradient which could disrupt the cloud on a dynamical timescale. Instead, the illuminated cloud must be in pressure equilibrium which also takes into account the radiation pressure (Różańska et al., 2006; Baskin et al., 2014). Some models also suggest the requirement of magnetic pressure for the stability of BLR clouds (e.g. Krause et al., 2012). Overall, the bright face has a much lower density, and the density is rising towards the unilluminated side. Clouds in pressure equilibrium were considered in several papers. Some of these computations were done in the context of the warm absorber (Różańska et al., 2006; Adhikari et al., 2015, 2019) but this medium may be the same as BLR. The change of the density and temperature is not like a power law but rather like a rapid transition (see e.g. Fig. 1 of Baskin et al. 2014, or Fig. 3 of Adhikari et al. 2019), so the LOC model would not reproduce it well, and eventually, a discontinuous jump in density and ionizing photon flux might be a better approximation. Computations of clouds under constant pressure are possible within CLOUDY environment, and this could be done in the future. Constant pressure assumption will allow for an approximate co-existence of material at strongly different local densities which would be forced by the pressure balance and the temperature gradient, although this would never represent yet the more complicated mixture of conditions which could form when the dynamical aspect of cloud formation/destruction is taken into account.
The second problem is the heating mechanism. Already the early papers asked the question of whether the Fe II emission forms as a result of the irradiation by the central source or due to mechanical heating. For example, Joly (1981) argued that either mechanical heating is the dominating channel of production, or the clouds must be very optically thick, in which case the irradiation flux is thermalized and the mechanism is similar to the mechanical heating.
The possibility of measuring the Fe II delay in the optical and UV bands contradicts the major collisional origin of the emission, but some contributions cannot be ruled out. Mechanical heating was discussed by Verner et al. (2003). The UV/optical ratio obtained in their paper is similar to what we measured from RM102 in the case of collisional heating, but radiative heating gave much higher values, 8 - 11, depending on the number of transitions. These values were obtained under a microturbulence velocity of 1 km s-1. They noted that Fe II optical emission flux is more sensitive to abundance than the Fe II (UV) band. Conversely, the Fe II (UV) band is more sensitive to microturbulence than the Fe II optical band. Also Baldwin et al. (2004) considered an option of mechanical heating but they overall favoured the radiative heating model with strong microturbulence. Such a model, however, requires two separate emitting regions for optical and UV, as mentioned above. Therefore, the combination of mechanical and radiative heating might be an interesting motion for future studies.